OpenStudy (anonymous):

Mandy bought a bag of peanuts to share with her friends. Trisha received 1/2 of the peanuts in the bag. Vince received 1/4 of the peanuts Trisha received. Ray received 1/3 of the peanuts Vince received. If Ray received 4 peanuts, how many peanuts were in the bag Mandy bought?

OpenStudy (anonymous):

12 maybe

OpenStudy (anonymous):

no its 96. i figured it out

OpenStudy (anonymous):

how is it 96?

OpenStudy (anonymous):

theres 2/10 of them left

OpenStudy (anonymous):

because 4 is 1/3 of 12, and 12 is 1/4 of 48 and 48 is 1/2 of 96

OpenStudy (anonymous):

true, i ddint read the rest! 420!

OpenStudy (anonymous):

B=Bag Trish = 1/2B Vince = 1/4trish or 1/4(1/2B) Ray = 1/3 Vince or 1/3 (1/4 (1/2B))). So if Ray is 4 you just start solving backwards 4*3 = 12 (that's what vince got 12 * 4 = 48 (that's what trish got 48 * 2 = bag.